A jumper (mass = 70 kg) lands with a peak force of 19 times body weight. If the load is concentrated on the facets, which have an area of 1 cm2, will they break? (Bone Ultimate Stress = 190 MPa, g = 9.8 m/s2). Hint: Remember to convert the cm to m twice since it is squared.

Yes

No
Answer #1

here,

Force = 19mg

area, A = 1 cm^2 = 1 *10^-4 m^2

here, for the maximum stress acting

maximum stress = force/area

maximum stress = 19 * 70 * 9.8/(1 *10^-4)

maximum stress = 1.303 *10^8 N/m^2 = 130.3 MPa

the stress acting is 130.3 MPa

hence, this is less than the ultimate stress of 190 MPa

hence ,NO the bone will NOT break